# Joint Discriminative-Generative Modeling via Dual Adversarial Training

### v1

@@ Abstract
Simultaneously achieving robust classification and high-fidelity generative modeling within a single framework presents a significant challenge. Hybrid approaches, such as Joint Energy-Based Models (JEM), interpret classifiers as EBMs but are often limited by the instability and poor sample quality inherent in SGLD-based training. We address these limitations by proposing a novel training framework that integrates adversarial training (AT) principles for both discriminative robustness and stable generative learning. The proposed method introduces three key innovations: (1) the replacement of SGLD-based JEM learning with a stable, AT-based approach that optimizes the energy function by discriminating between real data and PGD-generated contrastive samples using the BCE loss; (2) synergistic adversarial training for the discriminative component that enhances classification robustness while eliminating the need for explicit gradient penalties; and (3) a two-stage training procedure to resolve the incompatibility between batch normalization and EBM training. Experiments on CIFAR-10, CIFAR-100, and ImageNet demonstrate that our method substantially improves adversarial robustness over existing hybrid models while maintaining competitive generative performance. On ImageNet, when optimized for generative modeling, our model’s generative fidelity surpasses that of BigGAN and approaches diffusion models, representing the first MCMC-based EBM approach to achieve high-quality generation on complex, high-resolution datasets. Our approach addresses key stability issues that have limited JEM scaling and demonstrates that adversarial training can serve as an effective foundation for unified frameworks capable of generating and robustly classifying visual data. Code is available at https://github.com/xuwangyin/DAT.
@@ 1 Introduction
Deep learning models have traditionally been developed with either discriminative or generative objectives in mind, rarely excelling at both simultaneously. Discriminative models are optimized for classification or regression tasks but lack the ability to model data distributions, while generative models can synthesize new data samples but may underperform on downstream classification tasks. Recent research has explored unifying these approaches through joint discriminative-generative modeling frameworks that aim to combine the predictive power of discriminative approaches with the rich data understanding of generative models.
@@
    A stable AT-based alternative to SGLD-based JEM learning. We replace the unstable SGLD-based JEM learning with an adversarial training approach that optimizes the energy function through Binary Cross-Entropy loss using PGD-generated contrastive samples. This fundamentally addresses the training instabilities that have plagued JEM, enabling reliable convergence and significantly improved sample quality.
@@
    Two-stage training for batch normalization compatibility. We introduce a two-stage training strategy that reconciles the conflicting requirements of batch normalization for discriminative training and its incompatibility with EBM sampling, eliminating the need for alternative normalization techniques.
Experiments across datasets of increasing complexity, from CIFAR-10 to ImageNet, demonstrate that our approach scales effectively. Compared to existing hybrid models, out approach achieves substantially improved adversarial robustness while maintaining competitive generative performance. This unique combination of strong robustness and generative capabilities enables higher-quality counterfactual explanations—our model creates examples substantially more faithful to target class characteristics compared to non-robust and robustness-only methods. Furthermore, on ImageNet, when optimized for generative modeling, our model’s generative fidelity surpasses that of strong, specialized models like BigGAN and approaches the quality of diffusion models, demonstrating that EBM-based approaches can compete with other generative models. These results highlight the flexibility of our framework, demonstrating its capacity to function as both a state-of-the-art hybrid model and a powerful, standalone generative model with potential applications in a broader range of image synthesis tasks (Santurkar et al., 2019).
@@ 2 Related work
Joint discriminative-generative modeling The pursuit of joint discriminative-generative modeling, or hybrid modeling, aims to combine the predictive power of discriminative approaches with the rich data understanding of generative models within a single framework. This line of research is motivated by the potential to improve classifier robustness, calibration, and out-of-distribution detection, while also enabling tasks like sample generation (e.g., for counterfactual explanation) and semi-supervised learning. A significant thrust in this area involves Energy-Based Models (EBMs). Early work by Xie et al. (2016) showed how generative ConvNets could be derived from discriminative ones, framing them as EBMs. Du and Mordatch (2019) demonstrated that implicitly generative EBMs can achieve strong performance on discriminative tasks like adversarially robust classification and out-of-distribution detection, while addressing scalable EBM training challenges. Grathwohl et al. (2019) introduced Joint Energy-Based Models (JEM), which explicitly reinterpret standard classifiers as EBMs over the joint distribution of data and labels $p(x,y)$, allowing simultaneous classification and generation. Yang et al. (2023) incorporated sharpness-aware minimization (SAM) to smooth energy landscapes and removed data augmentation from the EBM loss term to improve both classification accuracy and generation quality of JEM. Guo et al. (2023) proposed EGC, which employs Fisher divergence within a diffusion framework to learn an unconditional score function $\nabla\log p(x)$ and a conditional classifier $p(y|x)$ for unified classification and generation, thereby circumventing the computational challenges of traditional energy-based model training.
@@
Our work builds upon these foundations by incorporating adversarial training principles into the joint modeling framework. We draw particularly from recent advances in adversarial training for EBMs (Yin et al., 2022) and methods for achieving robustness on both in-distribution and out-of-distribution data (Augustin et al., 2020). The connection between robust classifiers and generative capabilities has also been explored, with Santurkar et al. (2019) demonstrating that robust classifiers can perform various image synthesis tasks through gradient-based optimization. For a comprehensive discussion of these approaches, see Appendix A.1.
@@ 3.1 Joint Energy-Based Model
where $Z(\theta)$ is an intractable global normalizing constant. By marginalizing out the label $y$, a marginal density over the input data $x$ can be obtained:
@@
A JEM is trained by maximizing the joint log-likelihood $\log p_{\theta}(x,y)$ over labeled training datapoints $(x,y)$ drawn from an empirical joint distribution $p_{\text{data}}(x,y)$. The joint log-likelihood is typically factorized as $\log p_{\theta}(y|x)+\log p_{\theta}(x)$. The conditional term $\log p_{\theta}(y|x)$ can be maximized by minimizing the standard cross-entropy classification loss. The marginal term $\log p_{\theta}(x)$ is optimized using the EBM gradient:
@@ 3.2 Learning JEM with adversarial training
where $\alpha(x)=1-\sigma(-E_{\theta}(x))$ and $\beta(x)=\sigma(-E_{\theta}(x))$ are data-dependent scaling factors, and $\sigma$ denotes the logistic sigmoid function. This formulation preserves the structural form of Equation˜5 while introducing adaptive scaling factors that modulate gradient contributions according to the model’s current energy assignments. According to Yin et al. (2022), these scaling factors stabilize training by providing automatic gradient regularization: as $E_{\theta}(x)$ increases for $p_{\text{data}}$ samples, the corresponding scaling factor $\alpha(x)=1-\sigma(-E_{\theta}(x))$ approaches zero, thereby attenuating the gradient contribution from such samples and preventing numerical overflow; conversely, when $E_{\theta}(x)$ becomes very negative for contrastive samples, $\beta(x)=\sigma(-E_{\theta}(x))$ approaches zero, preventing numerical underflow. In contrast, the standard EBM gradient (Equation˜5) is unconstrained and permits $E_{\theta}(x)$ to achieve arbitrarily large or small magnitudes, resulting in numerical instability during optimization. This gradient formulation stabilizes training at the cost of limiting the EBM to modeling the support of $p_{\text{data}}$ rather than learning the full density.
@@
Minimizing this $\mathcal{L}_{\text{BCE}}$ implicitly trains the energy function $E_{\theta}(x)$ to assign low energy to data samples from $p_{\text{data}}(x)$ and high energy to the contrastive samples computed using the PGD attack.
@@ 3.3 Classifier robustness and implicit regularization
Implicit regularization. Incorporating AT for the classifier not only ensures robust accuracy but also yields a synergistic benefit for the generative component. Specifically, we find empirically that it eliminates the need for the $R_{1}$ gradient penalty (Mescheder et al., 2018), an explicit regularization required by the original AT-EBMs framework (Yin et al., 2022) that can constrain model expressiveness. Our analysis reveals that AT provides implicit regularization that encompasses $R_{1}$-style penalties through its first-order penalty on gradient norms (see Section˜A.2). While these two approaches operate at different scales and through different objectives, our experiments suggest that the local smoothness induced by AT provides sufficient implicit regularization for stable EBM training, reducing the need for explicit gradient penalties.
@@ 3.5 Two-stage training
A fundamental challenge in training joint models is the use of batch normalization (BN) (Ioffe and Szegedy, 2015). While BN is highly beneficial for stabilizing and speeding up standard deep network training, it is often found to interfere with the learning dynamics of EBMs and their sampling procedures (Grathwohl et al., 2019; Yin et al., 2022; Zhao et al., 2020, 2016). Consistent with these findings, we observe that enabling BN during joint training destabilizes the optimization of the generative modeling term $\mathcal{L}_{\text{BCE}}$, leading to oscillating losses and failure to converge. This instability arises from BN’s reliance on batch-dependent normalization statistics: as negative samples (e.g., from Equation˜6) change throughout training, the BN statistics $S_{t}$ continuously drift, effectively transforming the energy function from a fixed parameterization $E_{\theta}(x)$ into a time-varying one $E_{\theta,S_{t}}(x)$. In our experiments, we find that disabling BN removes this dependence and restores stable optimization of $\mathcal{L}_{\text{BCE}}$, confirming that the failure mode arises from computing BN statistics on the evolving negative samples.
However, simply disabling BN from the start would negatively impact the initial training of the robust classifier backbone. To reconcile these conflicting requirements, we implement a two-stage training strategy:
  • Stage 1 discriminative training. In this initial stage, we train the network with BN enabled, optimizing only the robust classification objective $\mathcal{L}_{\text{AT-CE}}$ (Equation˜11). This stage is equivalent to standard adversarial training and leverages BN to achieve faster convergence and strong robust classification performance. Notably, this stage can be skipped by initializing from pretrained off-the-shelf robust classifiers, making our approach immediately applicable to existing robust models.
  • Stage 2 joint training. After robust discriminative training, we disable BN throughout the network by setting the BN modules in eval mode. We then continue training by optimizing the complete objective function $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$ (Equation˜12).
While alternative approaches such as spectral normalization and virtual batch normalization have been considered for stabilizing EBM training (Zhao et al., 2020; Miyato et al., 2018; Zhao et al., 2016), our experiments demonstrate that this two-stage approach effectively addresses the BN incompatibility without requiring such alternatives. Disabling BN in Stage 2 enables stable generative loss convergence and dramatically improves generative performance, with minimal impact on the robust accuracy established in Stage 1 (see Section˜A.12 for training curves).
@@ 3.6 Data augmentation
Strong data augmentations are necessary for achieving robust classification (Rebuffi et al., 2021; Gowal et al., 2020) but can distort the data distribution in ways detrimental to generative modeling. We therefore follow Yang et al. (2023) and apply separate augmentation strategies to the discriminative and generative components. While Yang et al. (2023) concludes that augmentations like random cropping with padding should be excluded from generative training to avoid artifacts like black borders, we find this is not a limitation in our framework. Notably, even with random cropping and padding applied, our generated samples do not inherit these artifacts, allowing us to improve robustness without degrading sample quality (see Section˜A.4.2). We therefore apply strong augmentations to the discriminative term $\mathcal{L}_{\text{AT-CE}}$ and mild augmentations to the generative term $\mathcal{L}_{\text{BCE}}$.
@@ 4.1 Training setup
Datasets and architectures. We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009). For CIFAR-10/100 experiments, we use WRN-34-10 (Zagoruyko and Komodakis, 2016) following the official RATIO implementation. For ImageNet experiments, we employ ResNet-50 (He et al., 2016) and WRN-50-4 (Zagoruyko and Komodakis, 2016).
Two-stage training. Since Stage 1 training is equivalent to standard adversarial training, we use pretrained models when available: we use a pretrained CIFAR-10 model from RATIO (Augustin et al., 2020) and pretrained ImageNet models from Salman et al. (2020), while training our own CIFAR-100 model following Augustin et al. (2020). For Stage 2 training, we initialize from the Stage 1 model and continue joint training by setting the BN modules to eval mode (which disables BN while preserving the BN statistics computed during Stage 1). Complete training hyperparameters can be found in Section˜A.4.1.
Data augmentation. We employ separate data augmentation strategies for Stage 2 training: strong augmentations for $\mathcal{L}_{\text{AT-CE}}$ and basic transformations for $\mathcal{L}_{\text{BCE}}$ to preserve the data distribution. Detailed specifications can be found in Section˜A.4.2.
Out-of-distribution data. Same as RATIO, we use the 80 million tiny images (Torralba et al., 2008) as the OOD dataset ($p_{\text{ood}}$) for CIFAR-10/100 experiments. For ImageNet, as there are no established OOD datasets, we follow OpenImage-O (Wang et al., 2022) and construct an OOD dataset from Open Images training set (Krasin et al., 2016). We randomly sample 350K images, restricting our selection to those whose labels do not overlap with any ImageNet classes, yielding 300K samples for training and 50K for FID evaluation.

### v3

@@ Abstract
Simultaneously achieving robust classification and high-fidelity generative modeling within a single framework presents a significant challenge. Hybrid approaches, such as Joint Energy-Based Models (JEM), interpret classifiers as EBMs but are often limited by the instability and poor sample quality inherent in Stochastic Gradient Langevin Dynamics (SGLD)-based training. We address these limitations by proposing a novel training framework that integrates adversarial training (AT) principles for both discriminative robustness and stable generative learning. The proposed method introduces three key innovations: (1) the replacement of SGLD-based JEM learning with a stable, AT-based approach that optimizes the energy function by discriminating between real data and Projected Gradient Descent (PGD)-generated contrastive samples using the BCE loss; (2) synergistic adversarial training for the discriminative component that enhances classification robustness while eliminating the need for explicit gradient penalties; and (3) a two-stage training strategy that addresses normalization-related instabilities and enables leveraging pretrained robust classifiers, generalizing effectively across diverse architectures. Experiments on CIFAR-10/100 and ImageNet demonstrate that our approach: (1) is the first EBM-based hybrid to scale to high-resolution datasets with high training stability, simultaneously achieving state-of-the-art discriminative and generative performance on ImageNet $\times256$; (2) uniquely combines generative quality with adversarial robustness, enabling critical applications like robust counterfactual explanations; and (3) functions as a competitive standalone generative model, matching the generative quality of autoregressive methods (VAR-d16) and surpassing diffusion models while offering unique versatility.
@@ 1 Introduction
Deep learning models have traditionally been developed with either discriminative or generative objectives in mind, rarely excelling at both simultaneously (Ng and Jordan, 2001; Jebara, 2004; Lasserre et al., 2006; Xie et al., 2016; Grathwohl et al., 2019). Discriminative models are optimized for classification or regression tasks but lack the ability to model data distributions, while generative models can synthesize new data samples but may underperform on downstream classification tasks (Ng and Jordan, 2001; Jebara, 2004). Recent research has explored unifying these approaches through joint discriminative-generative modeling frameworks that aim to combine the predictive power of discriminative approaches with the rich data understanding of generative models (Xie et al., 2016; Lazarow et al., 2017; Jin et al., 2017; Du and Mordatch, 2019; Grathwohl et al., 2019; Chen et al., 2019; Guo et al., 2023; Deja et al., 2023).
@@
    A stable AT-based alternative to SGLD-based JEM learning. We replace the unstable SGLD-based JEM learning with an adversarial training approach that optimizes the energy function through Binary Cross-Entropy loss using Projected Gradient Descent (PGD; Madry et al. 2017)-generated contrastive samples. This fundamentally addresses the training instabilities that have plagued JEM, enabling reliable convergence and significantly improved sample quality.
@@
    Two-stage training strategy. We introduce a two-stage training strategy that provides practical benefits including leveraging pretrained robust classifiers and addressing normalization-related instabilities. This strategy generalizes across diverse architectures, working effectively for both ResNet (with batch normalization) and ConvNeXt (with layer normalization).
Experiments on CIFAR-10/100 and ImageNet demonstrate the effectiveness and scalability of our approach. Notably, our approach represents the first hybrid model to achieve competitive generative quality, strong classification performance, and adversarial robustness simultaneously—establishing that hybrid models need not compromise on any dimension. This establishes three advances for hybrid modeling:
  1. 1.
    First EBM-based hybrid that scales to high-resolution complex datasets. Prior EBM-based hybrid approaches (IGEBM, JEM, SADA-JEM) could not scale beyond low resolution or achieve competitive generative performance on ImageNet-level datasets due to SGLD instability. Our approach is the first to overcome these limitations, achieving competitive generative quality and strong classification performance on ImageNet $\times256$ with high training stability, demonstrating that EBM-based hybrid models can scale reliably to complex, high-resolution datasets.
  2. 2.
    Generative quality and adversarial robustness for critical applications. While other scalable hybrids like the diffusion-based EGC achieve reasonable performance, they do not reach state-of-the-art generative quality and lack adversarial robustness. Our dual capability of generative quality and adversarial robustness enables robust counterfactual explanations that are substantially more faithful to target class characteristics than non-robust or robustness-only methods (Section˜4.3.2), essential for interpretable ML in high-stakes domains.
  3. 3.
    Flexibility as a competitive generative model. When evaluated purely on generation quality, our approach (with ConvNeXt-Large) matches the autoregressive model VAR-d16 and surpasses diffusion models (ADM, LDM) on ImageNet $\times256$, while achieving significantly higher throughput than diffusion models. Beyond this competitive quality, our energy-based approach offers unique versatility for diverse synthesis tasks (inpainting, super-resolution, image manipulation) and compositional generation (Du et al., 2020; Santurkar et al., 2019).
These results demonstrate that hybrid models are no longer a compromise in either generative quality or classification performance—they can match state-of-the-art specialized models in both dimensions while providing the unique capabilities described above.
@@ 2 Related work
Joint discriminative-generative modeling The pursuit of joint discriminative-generative modeling, or hybrid modeling, aims to combine the predictive power of discriminative approaches with the rich data understanding of generative models within a single framework. This line of research is motivated by the potential to improve classifier robustness, calibration, and out-of-distribution detection (Grathwohl et al., 2019; Du and Mordatch, 2019), while also enabling tasks like sample generation (e.g., for counterfactual explanation (Deja et al., 2023)) and semi-supervised learning (Kingma et al., 2014). A significant thrust in this area involves Energy-Based Models (EBMs). Early work by Xie et al. (2016) showed how generative ConvNets could be derived from discriminative ones, framing them as EBMs. Du and Mordatch (2019) demonstrated that implicitly generative EBMs can achieve strong performance on discriminative tasks like adversarially robust classification and out-of-distribution detection, while addressing scalable EBM training challenges. Grathwohl et al. (2019) introduced Joint Energy-Based Models (JEM), which explicitly reinterpret standard classifiers as EBMs over the joint distribution of data and labels $p(x,y)$, allowing simultaneous classification and generation. Yang et al. (2023) incorporated sharpness-aware minimization (SAM) to smooth energy landscapes and removed data augmentation from the EBM loss term to improve both classification accuracy and generation quality of JEM. Guo et al. (2023) proposed EGC, which employs Fisher divergence within a diffusion framework to learn an unconditional score function $\nabla\log p(x)$ and a conditional classifier $p(y|x)$ for unified classification and generation, thereby circumventing the computational challenges of traditional energy-based model training.
@@
Our work builds upon these foundations by incorporating adversarial training principles into the joint modeling framework. Our approach is motivated by the theoretical insight that adversarially robust classifiers implicitly learn energy functions (Zhu et al., 2021; Wang et al., 2022b; Mirza et al., 2024), and we draw particularly from recent advances in adversarial training for EBMs (Yin et al., 2022) and methods for achieving robustness on both in-distribution and out-of-distribution data (Augustin et al., 2020; Korst and Asadulaev, 2022). Applications of robust classifiers to image synthesis are explored by Santurkar et al. (2019), who demonstrate that robust classifiers can perform various image synthesis tasks through gradient-based optimization. For a comprehensive discussion of these approaches, see Appendix A.1.
@@ 3.1 Joint Energy-Based Model
where $Z(\theta)=\sum_{y^{\prime}}\int\exp(f_{\theta}(x^{\prime})[y^{\prime}])dx^{\prime}$ is the partition function (an intractable global normalizing constant). By marginalizing out the label $y$, a marginal density over the input data $x$ can be obtained:
@@
A JEM is trained by maximizing the joint log-likelihood $\log p_{\theta}(x,y)$ over labeled training datapoints $(x,y)$ drawn from an empirical joint distribution $p_{\text{data}}(x,y)$. The joint log-likelihood is typically factorized as $\log p_{\theta}(y|x)+\log p_{\theta}(x)$. The conditional term $\log p_{\theta}(y|x)$ can be maximized by minimizing the standard cross-entropy classification loss. The marginal term $\log p_{\theta}(x)$ is optimized using the EBM gradient (LeCun et al., 2006):
@@ 3.2 Learning JEM with adversarial training
where $\alpha(x)=1-\sigma(-E_{\theta}(x))$ and $\beta(x)=\sigma(-E_{\theta}(x))$ are data-dependent scaling factors, and $\sigma$ denotes the logistic sigmoid function. This formulation preserves the structural form of Equation˜5 while introducing adaptive scaling factors that modulate gradient contributions according to the model’s current energy assignments. According to Yin et al. (2022), these scaling factors stabilize training by providing automatic gradient regularization: as $E_{\theta}(x)$ increases for $p_{\text{data}}$ samples, the corresponding scaling factor $\alpha(x)=1-\sigma(-E_{\theta}(x))$ approaches zero, thereby attenuating the gradient contribution from such samples and preventing numerical overflow; conversely, when $E_{\theta}(x)$ becomes very negative for contrastive samples, $\beta(x)=\sigma(-E_{\theta}(x))$ approaches zero, preventing numerical underflow. In contrast, the standard EBM gradient (Equation˜5) is unconstrained and permits $E_{\theta}(x)$ to achieve arbitrarily large or small magnitudes, resulting in numerical instability during optimization. This gradient formulation stabilizes training at the cost of limiting the EBM to modeling the support of $p_{\text{data}}$ rather than learning the full density. We provide a formal characterization of the learned distribution in Section A.2, where we show that the optimal solution under our joint objective learns $f_{\theta}^{*}(x)[y]=\log p_{\text{data}}(y|x)$ on the support with constant marginal energy $E_{\theta}^{*}(x)=0$.
@@
(i.e., $\nabla_{\theta}\mathcal{L}_{\text{BCE}}$ equals the right-hand side of Equation˜7). Minimizing this $\mathcal{L}_{\text{BCE}}$ implicitly trains the energy function $E_{\theta}(x)$ to assign low energy to data samples from $p_{\text{data}}(x)$ and high energy to the contrastive samples computed using the PGD attack.
@@ 3.3 Classifier robustness and implicit regularization
Implicit regularization. Incorporating AT for the classifier not only ensures robust accuracy but also yields a synergistic benefit for the generative component. The original AT-EBMs framework (Yin et al., 2022) required explicit $R_{1}$ gradient penalties (Mescheder et al., 2018) for training stability. However, explicit gradient penalties directly constrain the model and introduce an additional hyperparameter to tune, while AT achieves similar regularization implicitly through adversarial data augmentation, potentially preserving more model expressiveness. Building on Roth et al. (2020), we show that AT implicitly bounds the $R_{1}$ penalty (Section A.3), eliminating the need for explicit gradient regularization. We empirically validate this: AT maintains bounded $R_{1}$ gradients throughout training, while standard training exhibits gradient explosion (Figure 2). This confirms that AT provides sufficient implicit regularization for stable EBM training without explicit gradient penalties.
@@ 3.5 Two-stage training
Neural network architectures typically incorporate normalization layers to stabilize and speed up training: ResNet (He et al., 2016) uses batch normalization (BN) (Ioffe and Szegedy, 2015), while modern architectures like ConvNeXt (Liu et al., 2022) and Vision Transformers (Dosovitskiy et al., 2021; Vaswani et al., 2017) utilize layer normalization (Ba et al., 2016). Training energy-based joint models presents unique challenges with normalization layers. In particular, batch normalization has been identified as problematic for EBM training (Grathwohl et al., 2019; Yin et al., 2022; Zhao et al., 2020). Consistent with these findings, we observe that enabling BN during joint training destabilizes the optimization of the generative modeling term $\mathcal{L}_{\text{BCE}}$, leading to oscillating losses and failure to converge.
To address these challenges while maintaining the benefits of normalization during discriminative training, we propose a two-stage training strategy that generalizes effectively across diverse architectures:
  • Stage 1: Discriminative training. We first train the network with its original normalization configuration, optimizing only the robust classification objective $\mathcal{L}_{\text{AT-CE}}$ (Equation˜11). This stage is equivalent to standard adversarial training and leverages normalization layers to achieve faster convergence and strong robust classification performance. Notably, this stage can be skipped when pretrained robust classifiers are available, making our approach immediately applicable to existing robust models.
  • Stage 2: Joint training. After robust discriminative training, we modify the normalization behavior when necessary and continue training with the complete objective $\mathcal{L}(\theta)=\mathcal{L}_{\text{AT-CE}}(\theta)+\mathcal{L}_{\text{BCE}}(\theta)$ (Equation˜12). For architectures with batch normalization (ResNet, WRN), we disable BN by setting BN modules to eval mode, which freezes the BN statistics computed during Stage 1. For architectures with layer normalization (ConvNeXt), we maintain the normalization as-is.
This strategy not only addresses the incompatibility between batch normalization and EBM training, but also enables leveraging pretrained robust classifiers to dramatically reduce training costs (see Section˜B.2 for detailed computational analysis). As demonstrated in Section˜4.3, Stage 2 improves the generative modeling performance of pretrained robust classifiers while incurring minimal impact on the robust accuracy established in Stage 1 (see Section˜B.3 for training dynamics). Importantly, the two-stage training strategy works effectively for both ResNet and ConvNeXt models, making it suitable for architectures with better scaling properties such as Vision Transformers (Dosovitskiy et al., 2021; Singh et al., 2023; Peebles and Xie, 2023b).
@@ 3.6 Data augmentation
Strong data augmentations are necessary for achieving robust classification (Rebuffi et al., 2021; Gowal et al., 2020) but can distort the data distribution in ways detrimental to generative modeling. We therefore follow Yang et al. (2023) and apply separate augmentation strategies to the discriminative and generative components. While Yang et al. (2023) concludes that augmentations like random cropping with padding should be excluded from generative training to avoid artifacts like black borders, we find this is not a limitation in our framework. Notably, even with random cropping and padding applied, our generated samples do not inherit these artifacts, allowing us to improve robustness without degrading sample quality (see Section˜B.1.2). We therefore apply strong augmentations to the discriminative term $\mathcal{L}_{\text{AT-CE}}$ and mild augmentations to the generative term $\mathcal{L}_{\text{BCE}}$.
@@ 4.1 Training setup
Datasets and architectures. We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009). For CIFAR-10/100 experiments, we use WRN-34-10 (Zagoruyko and Komodakis, 2016) following the official RATIO implementation. For ImageNet experiments, we employ ResNet-50 (He et al., 2016), WRN-50-4 (Zagoruyko and Komodakis, 2016), and ConvNeXt-Large with ConvStem (Singh et al., 2023).
Two-stage training. Since Stage 1 training is equivalent to standard adversarial training, we use pretrained standard AT checkpoints when available: a standard AT checkpoint from the RATIO codebase (Augustin et al., 2020) for CIFAR-10, pretrained ImageNet ResNet-50 and WRN-50-4 models from Salman et al. (2020), and pretrained ConvNeXt-Large with ConvStem from Singh et al. (2023) (originally trained for $\ell_{\infty}=4/255$ robustness), while training our own CIFAR-100 model following Augustin et al. (2020). For Stage 2 training, we initialize from the Stage 1 model and continue joint training. For ResNet and WRN architectures, we set the BN modules to eval mode (which disables BN while preserving the BN statistics computed during Stage 1). Complete training hyperparameters can be found in Section˜B.1.1.
Data augmentation. We employ separate data augmentation strategies for Stage 2 training: strong augmentations for $\mathcal{L}_{\text{AT-CE}}$ and basic transformations for $\mathcal{L}_{\text{BCE}}$ to preserve the data distribution. Detailed specifications can be found in Section˜B.1.2.
Out-of-distribution data. Same as RATIO, we use the 80 million tiny images (Torralba et al., 2008) as the OOD dataset ($p_{\text{ood}}$) for CIFAR-10/100 experiments. For ImageNet, as there are no established OOD datasets, we follow OpenImage-O (Wang et al., 2022a) and construct an OOD dataset from Open Images training set (Krasin et al., 2016). We randomly sample 350K images, restricting our selection to those whose labels do not overlap with any ImageNet classes, yielding 300K samples for training and 50K for FID evaluation.